# DynaNoise: Dynamic Probabilistic Noise Injection for Defending Against Membership Inference Attacks

### v1

@@ 7.4. Experimental Setup and Results
Dataset Defense Model ($\uparrow$) Confidence ($\downarrow$) Loss ($\downarrow$) Shadow ($\downarrow$) $\text{MIDPUT}_{C}$ ($\uparrow$) $\text{MIDPUT}_{L}$ ($\uparrow$) $\text{MIDPUT}_{S}$ ($\uparrow$) $\text{MIDPUT}_{Overall}$ ($\uparrow$)
CIFAR10 None 0.8211 0.6956 0.7639 0.7841
SELENA (Tang et al., 2022) 0.7668 0.5446 0.6307 0.6755 0.0967 0.0789 0.0543 0.0766
DynaNoise 0.8156 0.2785 0.4221 0.5334 0.4116 0.3363 0.2452 0.331
ImageNet-10 None 0.9165 0.6489 0.7177 0.7612
SELENA (Tang et al., 2022) 0.7085 0.3893 0.5355 0.6817 0.0516 -0.0258 -0.1285 -0.0342
DynaNoise 0.9115 0.2248 0.308 0.5725 0.4191 0.4047 0.1837 0.3358
SST-2 None 0.8865 0.7913 0.7961 0.7854
SELENA (Tang et al., 2022) 0.8945 0.7774 0.7603 0.7415 0.0219 0.0438 0.0519 0.0392
DynaNoise 0.8911 0.2354 0.7863 0.735 0.5605 0.0144 0.0523 0.2091
Figures 3(a), 3(b), and 3(c) illustrate how DynaNoise responds to variations in base variance, lambda scale, and temperature on CIFAR-10. Increasing the base variance or lambda scale introduces more noise into high-risk predictions, which leads to a steady decline in membership inference attack success rates. Meanwhile, test accuracy remains relatively stable, indicating that low-sensitivity predictions are minimally affected. Temperature tuning further smooths the output distributions, effectively masking membership signals while preserving confidence in correct predictions. These observations validate the adaptive nature of DynaNoise in balancing noise and utility.
Figures 4(a), 4(b), and 4(c) show similar dynamics for ImageNet-10. As the base variance and lambda scale increase, noise adapts to prediction certainty, which selectively reduces overconfident outputs and lowers attack success. Adjusting the temperature parameter introduces a softening effect on logits, which initially helps reduce the distinguishability between member and non-member outputs. However, beyond a certain point, further increases in temperature yield diminishing returns, as the shadow attack success rate begins to rise again. This suggests that moderate temperature values are optimal for balancing privacy protection and output utility.
Figures 5(a), 5(b), and 5(c) display results for SST-2, where the target model already achieves high accuracy. Even with minimal impact on accuracy, tuning the defense parameters consistently lowers attack success rates, particularly with increased temperature. These results highlight that DynaNoise can still offer meaningful privacy gains in settings where the model is already well-optimized and has limited room for further improvement.
Moreover, across all datasets, as shown in Figure 3(c), 4(c), and 5(c), we observe that varying the temperature parameter has little to no effect on the test accuracy of the target model after DynaNoise is applied. This behavior stems from the role of temperature in post-processing the logits. To be more precise, it scales the output distribution without altering the predicted class. Since temperature is applied within the softmax function and does not change the relative ordering of logits in most cases, the top-1 prediction remains unaffected. Consequently, while higher temperatures effectively smooth the output probabilities and making it harder for an attacker to distinguish between member and non-member samples, they do so without degrading model performance. This characteristic further demonstrates the utility-preserving advantage of DynaNoise.
Overall, the figures confirm that DynaNoise’s parameterization enables fine-grained control over the privacy-utility trade-off, leveraging prediction sensitivity to selectively apply membership obfuscation where it matters the most.
(a) Base Variance Comparison
(b) Lambda Scale Comparison
(c) Temperature Comparison
Figure 3. CIFAR10: Accuracy and ASR comparison over varying (a) Base Variance, (b) Lambda Scale, and (c) Temperature.
(a) Base Variance Comparison
(b) Lambda Scale Comparison
(c) Temperature Comparison
Figure 4. ImageNet-10: Accuracy and ASR comparison over varying (a) Base Variance, (b) Lambda Scale, and (c) Temperature.
(a) Base Variance Comparison
(b) Lambda Scale Comparison
(c) Temperature Comparison
Figure 5. SST-2: Accuracy and ASR comparison over varying (a) Base Variance, (b) Lambda Scale, and (c) Temperature.
@@ 7.5. Time and Computational Overhead Analysis
The evaluated defense approaches differ not only in their per-sample time complexity but also in the overall computational resources required. DynaNoise perturbs the model’s logits by adding Gaussian noise whose variance is scaled based on query sensitivity, and then re-computes the $\ell$-dimensional probability vector using a softmax operation. This results in a per-sample computational overhead of
$\mathcal{O}(\ell),$
which corresponds to a single additional noise sampling and softmax computation. In practice, this represents a negligible runtime cost, making DynaNoise a lightweight and scalable defense. In contrast, SELENA employs an ensemble-based defense by training $K$ sub-models using the Split-AI architecture and subsequently applying self-distillation to consolidate their knowledge into a single model. During training, SELENA incurs a computational cost of
$\mathcal{O}(K\cdot C_{\text{model}}),$
where $C_{\text{model}}$ denotes the cost of training a single target model. While the self-distilled model reduces the inference cost to that of a standard target model, the training phase remains substantially more resource-intensive and time-consuming.
Beyond per-sample time complexity, the overall resource requirements are also critical. DynaNoise relies on a lightweight post-hoc noise injection approach, resulting in a memory footprint that remains nearly identical to that of the base model. In contrast, SELENA requires training and maintaining an ensemble of $K$ sub-models, which, even when parallelized, demands significantly more GPU memory and computational resources. As a result, DynaNoise provides a more efficient and scalable defense with substantially lower overall overhead.
7.6. Discussions and Limitations
Below we list the advantages and limitations of our proposed DynaNoise approach compared to SELENA and also in general.
@@ 7.6. Discussions and Limitations
  • Adaptive Noise Injection: DynaNoise dynamically adjusts the noise level based on query sensitivity. This adaptive mechanism enables DynaNoise to effectively mask membership signals while preserving high model accuracy.
  • Low Computational Overhead: Unlike SELENA, which requires training and maintaining an ensemble of sub-models, DynaNoise incurs only a minimal additional cost. Its lightweight post-hoc noise injection results in a resource footprint very similar to that of the target model.
  • Robust Membership Inference Defense Privacy–Utility Trade-off (MIDPUT): Our experiments demonstrate that DynaNoise consistently achieves higher MIDPUT values compared to SELENA. This indicates that the defense substantially reduces attack success rates with only a minor impact on model accuracy.
  • Simplicity and Scalability: DynaNoise’s implementation is straightforward and easily scalable to various datasets and model architectures without the complexities associated with ensemble training in SELENA.
@@
  • Hyperparameter Sensitivity: Similar to SELENA, the performance of DynaNoise depends on the tuning of key parameters (base variance, lambda scale, and temperature). Although extensive experiments have allowed us to identify robust parameter ranges, some degree of manual tuning remains necessary to optimize performance across different settings.
  • No Formal Privacy Guarantees: DynaNoise, like many practical defenses including SELENA, does not provide formal mathematical privacy guarantees. However, it is designed to offer strong empirical protection against membership inference attacks, as validated across diverse datasets, model architectures, and attack types.
@@ 8. Conclusion
In this paper, we proposed DynaNoise, an adaptive noise injection approach that dynamically adjusts the level of noise added to a model’s output based on the sensitivity of each query. In addition, we introduced a practical evaluation metric, called Membership Inference Defense Privacy–Utility Trade-off (MIDPUT), which measures how effectively a defense balances membership privacy gains with the associated cost in target model accuracy. Our extensive experimental analysis shows that DynaNoise consistently outperforms SELENA as the state-of-the-art baseline MIA defense approach. Unlike SELENA, which relies on an ensemble approach and incurs high computational overhead, DynaNoise operates with a lightweight post-hoc noise injection mechanism that demands minimal additional resources. This efficiency makes our approach scalable and well-suited for resource-constrained environments, while the adaptive noise modulation leads to a significant reduction in membership inference success rates. Future work will focus on applying such adaptive noise injection during the training phase. Additionally, we plan to explore the integration of DynaNoise with formal differential privacy techniques to provide stronger, provable privacy guarantees.

### v2

@@ 7.4. Experimental Setup and Results
Dataset Defense Model ($\uparrow$) Confidence ($\downarrow$) Loss ($\downarrow$) Shadow ($\downarrow$) LiRA ($\downarrow$) Entropy ($\downarrow$) M-Entropy ($\downarrow$)
CIFAR10 None 0.7875 0.6124 0.6180 0.6387 0.6097 0.6132 0.6386
AdvReg (Nasr et al., 2018) 0.7573 0.5516 0.5829 0.5708 0.5836 0.5482 0.5686
MemGuard (Jia et al., 2019) 0.7875 0.5948 0.6109 0.6122 0.5893 0.5814 0.6156
RelaxLoss (Chen et al., 2022) 0.7890 0.5555 0.5845 0.5665 0.5905 0.5433 0.5639
SELENA (Tang et al., 2022) 0.7674 0.5394 0.5173 0.5346 0.5164 0.5309 0.5326
HAMP (Chen and Pattabiraman, 2024) 0.7422 0.5000 0.5000 0.5000 0.5000 0.5000 0.5000
DynaNoise 0.7807 0.5014 0.5219 0.5053 0.5342 0.5016 0.5049
ImageNet-10 None 0.7958 0.5824 0.6003 0.5817 0.5721 0.5357 0.5855
AdvReg (Nasr et al., 2018) 0.7315 0.5244 0.5371 0.5426 0.5381 0.5312 0.5429
MemGuard (Jia et al., 2019) 0.7958 0.5800 0.6016 0.5817 0.5721 0.5333 0.5831
RelaxLoss (Chen et al., 2022) 0.7646 0.5203 0.5460 0.5416 0.5436 0.5196 0.5395
SELENA (Tang et al., 2022) 0.6804 0.4952 0.5010 0.5055 0.5034 0.5175 0.4993
HAMP (Chen and Pattabiraman, 2024) 0.6588 0.5000 0.5000 0.5000 0.5045 0.5000 0.5003
DynaNoise 0.7962 0.5316 0.5944 0.5549 0.5598 0.5139 0.5543
SST-2 None 0.8635 0.5162 0.5335 0.5571 0.5364 0.5558 0.5666
AdvReg (Nasr et al., 2018) 0.7092 0.5126 0.5037 0.5034 0.5154 0.5000 0.5000
MemGuard (Jia et al., 2019) 0.8635 0.5144 0.5335 0.5266 0.5364 0.5252 0.5311
RelaxLoss (Chen et al., 2022) 0.8394 0.5115 0.5125 0.5000 0.5147 0.5000 0.5000
SELENA (Tang et al., 2022) 0.8805 0.5270 0.5276 0.5511 0.5234 0.5351 0.5409
HAMP (Chen and Pattabiraman, 2024) 0.8900 0.5000 0.5000 0.5000 0.5000 0.5000 0.5000
DynaNoise 0.8612 0.5000 0.5014 0.5000 0.5023 0.5000 0.5000
Table 4. Proposed MIDPUT metrics on different datasets under various defense mechanisms.
Dataset Defense $\text{MIDPUT}_{Conf}(\uparrow)$ $\text{MIDPUT}_{Loss}(\uparrow)$ $\text{MIDPUT}_{Shadow}(\uparrow)$ $\text{MIDPUT}_{LiRA}(\uparrow)$ $\text{MIDPUT}_{Ent}(\uparrow)$ $\text{MIDPUT}_{M\text{-}Ent}(\uparrow)$ $\text{MIDPUT}_{Overall}(\uparrow)$
CIFAR10 AdvReg (Nasr et al., 2018) 0.0306 0.0049 0.0377 -0.0041 0.0348 0.0399 0.0240
MemGuard (Jia et al., 2019) 0.0176 0.0071 0.0086 -0.0001 0.0125 0.0110 0.0095
RelaxLoss (Chen et al., 2022) 0.0584 0.0350 0.0737 0.0207 0.0714 0.0762 0.0559
SELENA (Tang et al., 2022) 0.0529 0.0806 0.0840 0.0732 0.0623 0.0860 0.0732
HAMP (Chen and Pattabiraman, 2024) 0.0571 0.0627 0.0834 0.0544 0.0579 0.0833 0.0665
DynaNoise 0.1043 0.0893 0.1266 0.0687 0.1048 0.1270 0.1035
ImageNet-10 AdvReg (Nasr et al., 2018) -0.0062 -0.0010 -0.0251 -0.0302 -0.0598 -0.0216 -0.0240
MemGuard (Jia et al., 2019) 0.0024 -0.0014 0.0000 0.0000 0.0024 0.0024 0.0010
RelaxLoss (Chen et al., 2022) 0.0310 0.0231 0.0090 -0.0027 -0.0150 0.0149 0.0101
SELENA (Tang et al., 2022) -0.0282 -0.0161 -0.0391 -0.0467 -0.0972 -0.0292 -0.0428
HAMP (Chen and Pattabiraman, 2024) -0.0545 -0.0233 -0.0463 -0.0624 -0.0806 -0.0397 -0.0511
DynaNoise 0.0512 0.0062 0.0272 0.0127 0.0222 0.0316 0.0252
SST-2 AdvReg (Nasr et al., 2018) -0.1507 -0.1245 -0.1006 -0.1333 -0.0985 -0.0877 -0.1159
MemGuard (Jia et al., 2019) 0.0019 0.0000 0.0305 0.0000 0.0306 0.0354 0.0164
RelaxLoss (Chen et al., 2022) -0.0193 -0.0031 0.0332 -0.0023 0.0317 0.0425 0.0138
SELENA (Tang et al., 2022) 0.0097 0.0264 0.0265 0.0335 0.0412 0.0462 0.0306
HAMP (Chen and Pattabiraman, 2024) 0.0427 0.06 0.0836 0.0629 0.0823 0.0931 0.0708
DynaNoise 0.0139 0.0298 0.0548 0.0318 0.0535 0.0643 0.0414
Figures 3, 4, and 5 show how the three DynaNoise parameters, base variance, lambda scale, and temperature, affect target model accuracy and representative membership inference attacks across CIFAR-10, ImageNet-10, and SST-2. Each figure reports results before and after applying DynaNoise, enabling direct assessment of privacy–utility trade-offs under parameter variation. Varying the base variance (Figure 3) results in only minor changes in target model accuracy while yielding small but consistent reductions in attack success rates, particularly for LiRA and Shadow. The M-Entropy attack remains largely stable across this range, suggesting that base variance influences the overall scale of perturbation while preserving the underlying accuracy profile.
A similar trend is observed for the lambda scale (Figure 4), where increasing the parameter results in minimal changes to target model accuracy and modest, incremental reductions in attack success rates. This behavior indicates that the lambda scale primarily adjusts the magnitude of entropy-guided noise without substantially altering the overall privacy–utility profile. Temperature variation (Figure 5) exhibits a more pronounced effect, with higher temperatures consistently reducing attack success rates under LiRA, Shadow, and M-Entropy across all evaluated datasets while leaving accuracy largely unchanged. Taken together, these results suggest that lambda scaling and temperature smoothing influence privacy through different but complementary mechanisms, jointly shaping the privacy–utility behavior of DynaNoise.
(a) Target model
(b) LiRA
(c) Shadow
(d) M-Entropy
Figure 3. Base variance sweep. Target model accuracy and representative attack ASR (before vs. after DynaNoise) across CIFAR-10, ImageNet-10, and SST-2.
(a) Target model
(b) LiRA
(c) Shadow
(d) M-Entropy
Figure 4. Lambda scale sweep. Target model accuracy and representative attack ASR (before vs. after DynaNoise) across datasets.
(a) Target model
(b) LiRA
(c) Shadow
(d) M-Entropy
Figure 5. Temperature sweep. Target model accuracy and representative attack ASR (before vs. after DynaNoise) across datasets.
@@ 7.5. Time and Computational Overhead Analysis
The evaluated defense mechanisms differ substantially in their computational requirements during both training and inference. Empirical measurements of training and inference overhead are summarized in Appendix C. We next discuss the computational complexity of each defense.
DynaNoise perturbs model logits by adding Gaussian noise whose variance adapts to query sensitivity, followed by a softmax recomputation of the $\ell$-dimensional probability vector. This operation has per-sample cost $\mathcal{O}(\ell)$, corresponding to one noise draw and a single softmax evaluation. As a post-hoc defense, DynaNoise imposes minimal runtime overhead and requires no retraining, making it lightweight and easily deployable.
Adversarial Regularization introduces a min–max training framework in which an auxiliary membership attack model is jointly optimized with the target classifier. Each training step includes additional attacker updates, yielding an overall training complexity of $\mathcal{O}(C_{\text{model}}+k\cdot C_{\text{attack}})$, where $C_{\text{attack}}$ denotes the attack model update cost and $k$ the number of attack iterations per batch. As a result, this defense requires retraining the target model and increases the complexity of the training procedure, while inference cost remains identical to that of the base classifier.
RelaxLoss modifies the training objective by penalizing overly confident predictions via a confidence-regularization term. This defense requires no auxiliary models and introduces no stochastic components, resulting in training complexity $\mathcal{O}(C_{\text{model}})$ and unchanged inference cost. Thus, RelaxLoss retains training efficiency comparable to the corresponding undefended baseline.
MemGuard protects against score-based membership inference attacks by solving a constrained optimization problem at inference time to minimally perturb the model’s posterior distribution. For each query, MemGuard performs an inner-loop iterative optimization over the output logits, with complexity approximately $\mathcal{O}(T\cdot\ell)$, where $T$ is the number of gradient steps and $\ell$ is the number of classes. While model retraining is not required, inference-time overhead is significantly higher than DynaNoise, RelaxLoss, or AdvReg, since a full iterative procedure must be run for every input. This makes MemGuard substantially more computationally expensive and less suited for real-time or large-scale deployments.
SELENA trains an ensemble of $K$ sub-models under the Split-AI framework and performs self-distillation to aggregate their predictions. This results in a total training complexity of $\mathcal{O}(K\cdot C_{\text{model}})$ and a correspondingly large memory footprint. After distillation, inference cost becomes equivalent to a single model, but the training phase is considerably more resource-intensive. Consistent with this design, empirical inference latency after distillation matches the baseline model (Appendix C).
HAMP combines training-time entropy regularization with inference-time output modification. During training, HAMP enforces high-entropy soft labels and an entropy-maximization term in the loss, resulting in training complexity $\mathcal{O}(C_{\text{model}})$ with no auxiliary models. At inference time, HAMP generates random reference inputs and replaces the sorted output probabilities of the queried sample with those from the random inputs while preserving label ordering. This procedure introduces an additional forward pass per query and a sorting operation over the output distribution, yielding an inference-time complexity of approximately $\mathcal{O}(C_{\text{model}}+\ell\log\ell)$. While HAMP avoids retraining multiple models, its inference overhead is higher than that of DynaNoise due to the extra forward evaluation and probability replacement step, yielding approximately a $2\times$ inference-time overhead in practice (Appendix C).
Overall. DynaNoise offers the most computationally efficient defense among all evaluated methods. Unlike RelaxLoss, AdvReg, SELENA, and HAMP, it operates entirely post-hoc, requires no retraining, introduces no auxiliary models, and adds only an $\mathcal{O}(\ell)$ inference-time operation. Compared to MemGuard and HAMP, which incur additional per-query optimization or extra forward passes, DynaNoise remains substantially lighter at inference time. As a result, DynaNoise achieves strong privacy protection with the smallest computational and memory footprint, making it well suited for deployment on pretrained models in resource-constrained or large-scale settings.
7.6. Discussion and Limitations
DynaNoise provides a lightweight, post-hoc defense against membership inference attacks, achieving strong privacy protection with minimal impact on model accuracy and inference cost. We next discuss its practical advantages and limitations.
@@ 7.6. Discussion and Limitations
  • Distribution-aware sensitivity: Unlike SELENA, AdvReg, RelaxLoss, MemGuard, and HAMP, which apply fixed training-time regularization,ensemble aggregation, or global inference-time output modification, DynaNoise adapts its perturbation level using a continuous, entropy-based measure of prediction sensitivity. By leveraging the full predictive distribution on a per-query basis, DynaNoise selectively injects noise for high-risk predictions while avoiding unnecessary perturbations on low-risk queries.
  • Post-hoc and model-agnostic: The defense operates purely at inference time and requires no retraining, architectural changes, or access to training data. In contrast, AdvReg, RelaxLoss, and HAMP modify the training procedure, SELENA trains an ensemble of sub-models, and MemGuard solves a per-query constrained optimization problem over logits. DynaNoise can therefore be deployed immediately on pretrained models such as AlexNet or DistilBERT, making it well suited for MLaaS and black-box settings.
  • Efficiency and simplicity: DynaNoise introduces only three interpretable hyperparameters ($\sigma_{0}$, $\lambda$, and $T$) and adds a single noise sampling and softmax evaluation per query, resulting in $\mathcal{O}(\ell)$ inference-time overhead. In comparison, RelaxLoss and HAMP require retraining with modified objectives, SELENA incurs ensemble training costs, AdvReg relies on adversarial optimization, and MemGuard performs iterative per-query perturbation solving. As a result, DynaNoise remains the lightest-weight option, with the lowest computational and memory footprint among all evaluated defenses.
  • Superior privacy-utility trade-off: Across CIFAR-10 and ImageNet-10, DynaNoise achieves the highest overall MIDPUT values, delivering strong reductions in attack success rates while preserving near-baseline target model accuracy. On SST-2, DynaNoise provides competitive privacy gains with stable accuracy, closely matching the best-performing defenses. Compared to SELENA, AdvReg, RelaxLoss, MemGuard, and HAMP, DynaNoise consistently exhibits favorable privacy-utility behavior across diverse datasets and attack types.
@@
  • Hyperparameter tuning: Similar to SELENA, AdvReg, RelaxLoss, MemGuard, and HAMP, the effectiveness of DynaNoise depends on selecting appropriate hyperparameters, including its variance-scaling and temperature components. While we provide stable default ranges that perform well across datasets, some dataset-specific tuning may be required to achieve optimal privacy-utility trade-offs.
  • No formal privacy guarantees: Like SELENA, AdvReg, RelaxLoss, MemGuard, and HAMP, DynaNoise provides empirical robustness against membership inference attacks but does not offer formal differential privacy guarantees. Developing theoretical privacy bounds for entropy-aware, query-adaptive output perturbation remains an important direction for future work.
  • Vulnerability to repeated queries: As with other stochastic inference-time defenses, including MemGuard, HAMP, and DynaNoise, repeated or highly similar queries may allow an adversary to average out injected randomness and partially recover the underlying output distribution (Rahimian et al., 2020). DynaNoise therefore assumes a limited-query threat model; potential mitigations include enforcing per-record query limits and detecting suspicious query patterns using stateful or similarity-based monitors (Li et al., 2022, 2025).
  • Label-only membership inference: DynaNoise protects against score-based membership inference attacks and does not defend against label-only MIAs that rely solely on predicted labels. Extending defenses to this threat model is an important direction for future work.
@@ 8. Conclusion
In this work, we introduced DynaNoise, a lightweight and adaptive inference-time defense against membership inference attacks. Unlike prior approaches that require retraining, ensemble models, or uniform confidence suppression, DynaNoise operates as a purely post-hoc mechanism that dynamically injects calibrated noise based on query sensitivity. By modulating protection strength using entropy, DynaNoise selectively shields high-risk predictions while preserving the utility of low-risk queries. Extensive evaluation across vision and language benchmarks demonstrates that this query-adaptive strategy consistently reduces attack success rates while maintaining near-baseline accuracy, outperforming training-time, post-processing, and ensemble-based defenses in terms of overall privacy–utility trade-off.
To facilitate holistic evaluation of inference-time defenses, we also proposed the MIDPUT metric as a complementary summary measure capturing the balance between privacy protection and predictive performance. Across diverse datasets and attack settings, DynaNoise achieves strong MIDPUT values, highlighting the effectiveness of adapting protection to query sensitivity rather than applying global perturbations. These results indicate that query-aware noise injection offers a practical and resource-efficient path toward improving privacy in deployed machine learning systems. Future work will explore extending the DynaNoise framework with alternative sensitivity estimators, such as modified entropy, and integrating adaptive noise mechanisms with formal privacy techniques to further strengthen robustness while retaining strong empirical performance.
Appendix A Entropy–Sensitivity Correlation Plots
This section contains the full entropy–Jacobian sensitivity correlation plots, which provide empirical support for the analysis in Section 6.1, as illustrated in Figure 6.
(a) CIFAR-10 (AlexNet), r=0.41r=-0.41
(b) ImageNet-10 (AlexNet), r=0.46r=-0.46
(c) SST-2 (DistilBERT), $r=0.196$
Figure 6. Correlation between output entropy and Jacobian-based sensitivity across datasets and model architectures. These plots support the analysis in Section 6.1.
Appendix B Dataset-Level Confidence and Entropy Analysis
To better understand the dataset-dependent behavior of membership inference attacks (MIAs), we analyze the distributions of maximum prediction confidence and predictive entropy for training (IN) and non-training (OUT) samples across all evaluated datasets. These statistics correspond directly to the decision signals exploited by confidence-, entropy-, and modified-entropy-based MIAs, and provide insight into both the source of membership leakage and the mechanisms by which DynaNoise mitigates it. The resulting percentile plots for CIFAR-10, ImageNet-10, and SST-2 are shown in Figures 7, 8, and 9, respectively.
B.1. Methodology
For each dataset, we compute the maximum softmax probability and the Shannon entropy of the model’s predicted class distribution for every query. We report percentile curves for IN and OUT samples separately, both before applying any defense and after applying DynaNoise. These curves reveal how rapidly confident or low-entropy predictions emerge for training samples relative to non-members, and how this separability changes under adaptive noise injection.
B.2. CIFAR-10
On CIFAR-10, we observe a moderate but consistent separation between IN and OUT samples in both confidence and entropy. Training points tend to reach higher confidence and lower entropy earlier than non-members, although the gap increases gradually rather than abruptly. This behavior explains why metric-based MIAs achieve non-trivial but limited success on this dataset.
After applying DynaNoise, the confidence and entropy gaps between IN and OUT samples are noticeably compressed across the percentile range. Importantly, this compression occurs without collapsing the overall prediction structure, consistent with the minimal utility degradation observed in Table 3. These results illustrate that DynaNoise reduces excessive confidence disparities while preserving the underlying ranking of predictions.
Refer to captionRefer to caption
(a) Maximum confidence before (left) and after (right) DynaNoise
Refer to captionRefer to caption
(b) Predictive entropy before (left) and after (right) DynaNoise
Figure 7. CIFAR-10 percentile plots comparing confidence- and entropy-based statistics before and after applying DynaNoise.
B.3. ImageNet-10
ImageNet-10 exhibits the most pronounced IN/OUT separation among the evaluated datasets. Before applying any defense, training samples attain near-unit confidence and near-zero entropy at substantially lower percentiles than non-members, indicating a highly peaked prediction regime for memorized examples. This sharp divergence explains the strong performance of confidence-, entropy-, and LiRA-based MIAs on this dataset.
After applying DynaNoise, both confidence and entropy curves show a substantial reduction in IN/OUT separation. The early-percentile confidence spike for training samples is strongly suppressed, and entropy distributions for members and non-members become significantly more aligned across most of the percentile range. This directly explains the observed reduction in attack success rates under strong attacks and highlights the advantage of query-adaptive, entropy-scaled noise over global confidence suppression.
Refer to captionRefer to caption
(a) Maximum confidence before (left) and after (right) DynaNoise
Refer to captionRefer to caption
(b) Predictive entropy before (left) and after (right) DynaNoise
Figure 8. ImageNet-10 percentile plots comparing confidence- and entropy-based statistics before and after applying DynaNoise.
B.4. SST-2
In contrast to the vision datasets, SST-2 shows minimal separation between IN and OUT samples in both confidence and entropy even before applying defenses. Predictions remain relatively well-calibrated across percentiles, reflecting the behavior of pretrained transformer-based language models. Consequently, metric-based MIAs are intrinsically weaker on SST-2.
Applying DynaNoise results in only minor changes to the confidence and entropy distributions, further indicating that the method adapts naturally to low-risk regimes by avoiding unnecessary perturbations. This behavior aligns with the strong utility preservation observed for SST-2 while still providing protection against high-risk, low-entropy queries.
Refer to captionRefer to caption
(a) Maximum confidence before (left) and after (right) DynaNoise
Refer to captionRefer to caption
(b) Predictive entropy before (left) and after (right) DynaNoise
Figure 9. SST-2 percentile plots comparing confidence- and entropy-based statistics before and after applying DynaNoise.
B.5. Discussion
Across all datasets, these results demonstrate that membership leakage is fundamentally shaped by dataset complexity and model calibration. ImageNet-10’s low-entropy, high-confidence prediction regime creates favorable conditions for metric-based MIAs, whereas SST-2’s calibrated outputs inherently limit attack effectiveness. By scaling noise injection to per-query entropy, DynaNoise adapts naturally to these dataset-specific characteristics, reducing confidence-based separability where it is most pronounced while preserving utility in lower-risk regimes.
Appendix C Additional Computational Overhead Results
C.1. Empirical Training and Inference Overhead
Table 5 reports empirical measurements of training and inference overhead for all evaluated defenses. Inference latency is measured as the end-to-end per-sample runtime at deployment, including any defense-specific post-processing applied to model outputs. Inference overhead is reported relative to the undefended baseline model. Training overhead is reported as the ratio between the total training cost of a defended model and the corresponding undefended baseline.
Table 5. Empirical computational overhead of evaluated defenses. Inference latency is measured in milliseconds per sample. Training overhead is reported relative to the undefended baseline.
Defense End-to-End Inference Inference Overhead Training Overhead Retraining Required
(ms/sample) ($\times$ baseline) ($\times$ baseline)
None 0.1414 1.0 1.0 No
AdvReg 0.1414 1.0 1.0 Yes
MemGuard 28.1118 198.8 1.66 No
RelaxLoss 0.1414 1.0 1.0 Yes
SELENA 0.1414 1.0 2.61 Yes
HAMP 0.2819 2.0 1.0 No
DynaNoise 0.1432 1.01 1.0 No
C.2. Overhead Definitions
Let $T_{\text{model}}$ denote the per-sample forward-pass latency of the undefended target model, and let $T_{\text{defense}}$ denote any additional defense-specific computation performed at inference time. The end-to-end inference latency is defined as
(1) $T_{\text{e2e}}=\frac{1}{N}\sum_{i=1}^{N}\bigl(T_{\text{model}}^{(i)}+T_{\text{defense}}^{(i)}\bigr),$
where $N$ is the number of evaluated samples.
Inference overhead is reported as a multiplicative ratio relative to the undefended baseline:
(2) $\text{Inference Overhead}=\frac{T_{\text{e2e}}^{\text{defended}}}{T_{\text{e2e}}^{\text{baseline}}}.$
Training overhead is defined as the ratio between the total training cost of a defended model and that of the corresponding undefended baseline.
(3) $\text{Training Overhead}=\frac{T_{\text{train}}^{\text{defended}}}{T_{\text{train}}^{\text{baseline}}}.$
Here, $T_{\text{train}}^{\text{defended}}$ includes all training necessary for deployment, which may consist of retraining the target model (for training-time defenses) or training auxiliary components (e.g., defense classifiers), but does not include baseline models trained solely for experimental comparison.
C.3. Discussion
Post-hoc defenses exhibit markedly different computational profiles. MemGuard incurs extreme inference-time overhead due to per-query iterative optimization, resulting in orders-of-magnitude slower inference. HAMP approximately doubles inference latency due to an additional forward evaluation and output manipulation step. Training-time defenses such as AdvReg, RelaxLoss, and SELENA require retraining the target model; among these, SELENA exhibits substantially higher training cost due to training multiple sub-models and a subsequent distillation phase, while AdvReg and RelaxLoss have training costs comparable to the corresponding undefended baseline.
In contrast, DynaNoise operates entirely post hoc, requires no retraining or auxiliary models, and adds only a small ($\approx$1%) inference-time overhead corresponding to a single noise injection and softmax operation. As a result, DynaNoise achieves strong privacy protection with the smallest combined training and inference overhead among all evaluated defenses, making it particularly well suited for deployment on pretrained models in large-scale or resource-constrained settings.